If the empirical formula of a compound is P2O3, and it has a molecular weight equal to 220 g/mol, what is its molecular formula?


1
Expert's answer
2022-03-10T10:46:58-0500

Solution:

Empirical formula = P2O3

Molar mass of P2O3 = 2×Ar(P) + 3×Ar(O) = 2×31 + 3×16 = 110 (g/mol)

Molecular formula = (P2O3)n

Molar mass of (P2O3)n is 220 g/mol


n = (Molecular formula / Empirical formula) = (Molar mass of (P2O3)/ Molar mass of P2O3)

n = 220 / 110 = 2

Therefore,

Molecular formula = (P2O3)2 = P4O6


Answer: The molecular formula of a compound is P4O6